A nozzle is designed to expand air from 689 π‘˜π‘ƒπ‘Žπ‘Ž and 32℃ to 138 π‘˜π‘ƒπ‘Žπ‘Ž. Assume an isentropic expansion and negligible approach velocity. The airflow rate is 1.36 kg/s. Calculate (a) the exit velocity; (b) the proper exit cross-sectional area.


1
Expert's answer
2021-10-28T08:51:28-0400


(a) Since the expansion is isentropic

"\frac{T_2}{T_1} = (\frac{p_2}{p_1})^{γ-1/γ} \\

\frac{T_2}{305} = (\frac{138}{689})^{1.4-1/1.4} \\

T_2 = 192.625 \;K"

Applying steam flow energy equation to nozzle

"m(h_1 + \frac{v^2_1}{2000} + \frac{z_1g}{1000}) + Q = m(h_2 + \frac{v^2_2}{2000} + \frac{z_2g}{1000}) + w \\

m(h_1 + 0 +0) + 0 = m(h_2 + \frac{v^2_2}{2000} + 0) + 0"

Neglecting potential energy change

"mh_1 = m(h_2 + \frac{v^2_2}{2000}) \\

mc_pT_1 = m(c_pT_2 + \frac{v^2_2}{2000}) \\

1.36 \times 1.005 \times 305 = 1.36(1.005 \times 192.652 + \frac{v^2_2}{2000}) \\

v_2 = 475.205 \;m/s"

(b) Mass flow rate "m =\u03c1_1A_1v_1 = \u03c1_2A_2v_2"

From ideal gas equation

"p_2 = ρ_2RT_2 \\

ρ_2 = \frac{p_2}{RT_2} = \frac{138}{0.287 \times 192.652} = 2.4958 \;kg/m^3 \\

m=ρ_2A_2V_2 \\

1.36 = 2.4958 \times A_2 \times 475.205 \\

A_2 = 1.1467 \times 10^{-3} \;m^2"